# Rolling the DICE on Idiomaticity: How LLMs Fail to Grasp Context

### v1

@@ Abstract
Human processing of idioms relies on understanding the contextual sentences in which idioms occur, as well as language-intrinsic features such as frequency and speaker-intrinsic factors like familiarity. While LLMs have shown high performance on idiomaticity detection tasks, this success may be attributed to reasoning shortcuts in existing datasets. To this end, we construct a novel, controlled contrastive dataset designed to test whether LLMs can effectively use context to disambiguate idiomatic meaning. Additionally, we explore how collocational frequency and sentence probability influence model performance. Our findings reveal that LLMs often fail to resolve idiomaticity when it is required to attend to the surrounding context, and that models perform better on sentences that have higher likelihood. The collocational frequency of expressions also impacts performance. We make our code and dataset publicly available. [^1] https://github.com/mi-m1/dice
@@
Maggie Mi1 Aline Villavicencio1,2 Nafise Sadat Moosavi1 1University of Sheffield 2University of Exeter zmi1@sheffield.ac.uk a.villavicencio@exeter.ac.uk n.s.moosavi@sheffield.ac.uk
@@ 1 Introduction
Idiomatic expressions (IEs) are strange birds whose meaning may not be straightforwardly related to the meaning of the component words in isolation. For example, proficient English speakers understand “spill the beans” not as causing legumes to fall, but as disclosing a secret. Estimates suggest that there are 25,000 fixed expressions in English alone (Wèinreich, 1969), and a similar estimate is quoted for French (Gross, 1982). Notably, this figure is comparable to the order of magnitude of individual words in the lexicon (Jackendoff, 1997). This suggests that idioms are not mere linguistic curiosities but fundamental components of language.
Additionally, some of these are ambiguous, “potentially idiomatic expressions" (PIEs) that can be interpreted either non-compositionally (figuratively or idiomatically) or compositionally (literally), depending on the context in which they appear. Accurately identifying the meaning of a PIE often depends on its context and is essential for numerous downstream applications, such as machine translation (Dankers et al., 2022; Barreiro et al., 2013; Salton et al., 2014; Fadaee et al., 2018), sentiment analysis (Williams et al., 2015; Liu et al., 2017), and automatic spelling correction Horbach et al. (2016). Beyond these applications, it is also crucial to grasp the underlying meaning of the text.
Existing datasets that feature expressions with both literal and idiomatic usages often fail to rigorously examine the role of context in disambiguating meaning. This shortcoming arises because the literal meanings in such datasets typically result from syntactic modifications or semantic shifts, which inherently disrupt the idiomaticity of the expression. For example, passivization (2) or modification to the expression (3) often strip the expression of its idiomatic meaning, as shown by the idiomatic usage in (1). Consequently, this allows models to exploit surface-level differences, such as changes in grammatical structure or added modifiers, as shortcuts for determining literalness, rather than relying on their understanding of idiomaticity in context as the task intends. This undermines the evaluation’s effectiveness by encouraging models to use superficial cues instead of deeper contextual comprehension.
@@
[…]
In this paper, we address this gap by introducing a new benchmark designed to evaluate the of ability of LLMs to interpret idiomatic expressions in context. We focus on idioms, as these figurative expressions may serve as key indicators of a model’s linguistic understanding. Given the possibility of a predominant sense, our dataset presents idiomatic expressions, balancing both literal and figurative contexts, challenging models to rely on contextual cues for accurate interpretation. This contrastive evaluation forces models to distinguish between meanings based solely on context, preventing them from relying on memorized idioms. We hypothesize that if models do not depend on memorization, they will perform equally well in both senses of a PIE.
Moreover, in human language processing, frequency (Swinney and Cutler, 1979) and familiarity (Nippold and Rudzinski, 1993), alongside context (Estill and Kemper, 1982; Gibbs Jr et al., 1989; Gibbs and Nayak, 1989), are known to influence idiom comprehension and speed of processing (Tabossi et al., 2009). This may similarly apply to models. In addition to context, we analyze the impact of a language-intrinsic feature (expression frequency) and a model-intrinsic feature (likelihood) on model performance.
@@
(1) We present a comprehensive and robust evaluation dataset, containing PIEs no syntactical changes appearing in both figurative and literal settings. (2) Through fine-grained evaluations, we find that models struggle to use context for idiomaticity processing. (3) Based on frequency estimates, we find frequency is not a “free lunch” and that, whilst highly frequent idioms maybe be more likely to be disambiguated correctly, there is a trade-off between performance on the two idiomaticity settings (literal/figurative). (4) For models that learned idiomaticity sufficiently well, we observe that there is a strong relationship between the likelihood of the contextual sentence and performance on idiomaticity detection.
@@ 2 Related Works
Existing Datasets
@@
The biggest dataset for idiomatic sense disambiguation is MAGPIE (Haagsma et al., 2020). MAGPIE contains a total of 56,622 PIE instances, across 1,756 idioms. However, a large amount of deviation of the form of an expression was allowed when MAGPIE was curated. As a result, most of the literal uses of PIEs involve modifications to the form of the expression (similar to the example in Section 1). Other large datasets targeting various types of IEs have been released: The VNC-Tokens dataset focusing on verb-noun combinations Cook et al. (2008), IDIX on verb-noun phrase or verb-prepositional phrase expressions (Sporleder et al., 2010). SemEval-2013 has unrestricted expressions (Korkontzelos et al., 2013), and AStitchInLanguageModels contains noun compounds (Tayyar Madabushi et al., 2021).
@@
To address these problems: (1) we propose a novel evaluation set (DICE) where we strictly control the form of idiomatic expressions. This eliminates the possibility that models rely on grammatical variations for idiomaticity disambiguation. By maintaining consistent forms across contexts, our dataset ensures that the challenge lies in understanding contextual nuances, thereby providing a more accurate assessment of a model’s idiomatic comprehension. (2) Visibly, these datasets often only contain expressions of a singular type. As a result, we address this lack of coverage by including expressions that are both phrasal expressions and noun compounds. Moreover, we believe it is crucial to maintain the same form of expression in both literal and figurative contexts, so we aim to construct a more controlled evaluation framework.
Idiomaticity Detection
The task of idiomaticity sense disambiguation (ISD), or, idiomaticity detection involves evaluating whether an expression is used literally or figuratively in a sentence (Liu and Hwa, 2018; Salehi et al., 2014; Senaldi et al., 2016; Gharbieh et al., 2016). LLMs seem to achieve high performance in idiomaticity detection tasks (Phelps et al., 2024; Zeng and Bhat, 2021). This task can be framed as a binary classification, where models must determine whether the usage of a PIE is literal or idiomatic. However, their ability to accurately model context relevant for idiomatic understanding is unclear Garcia et al. (2021b) and has yet to be comprehensively determined. Given the crucial role of context in interpreting these expressions, it is important to evaluate to what extent models use context to make this distinction. However, since these models are trained on extensive corpora that likely include idiomatic expressions, it is unclear whether they are memorizing the surface forms or genuinely comprehending the context to detect the idiomatic usage.
Contrastive Evaluation
Contrastive evaluation often takes the form of minimal pairs evaluation, where a single perturbation such as a change in a word or condition is systematically introduced into otherwise identical conditions Prasad et al. (2019); Garcia et al. (2021a). This method has been noted for its advantage in identifying specific weaknesses in model understanding and robustness (Linzen et al., 2016; Sennrich, 2017; Robertson, 2019).
Our work fits within this category of contrastive evaluation, with a specific focus on context comprehension in idiomaticity detection. By presenting PIEs in both their figurative and literal forms, our dataset forces models to make use of subtle contextual cues to differentiate between meanings. Additionally, by controlling for the form of the expressions, we ensure they remain identical across contexts. Thus, the challenge stems from interpreting the surrounding context, rather than recognizing variations in form.
Memorisation and Context
@@
Pretrained language models tend to handle IEs mainly through memorization of stored expressions and token distributions rather than reasoning about the meaning in context. This reliance on memorization becomes evident when models face novel noun compounds. For instance, Li et al. (2022) found that while GPT-3’s interpretations of novel compounds closely matched human responses, its performance faltered when tasked with interpreting nonsensical strings. Unlike humans, who can use surrounding context to deduce meaning, GPT-3 failed to do so, indicating its heavy dependence on pre-learned token distributions rather than contextual clues. Similarly, Coil and Shwartz (2023) explored noun compound interpretation using GPT-3 and found that, although the model performs well with common noun compounds, its performance drops significantly with novel compounds. This suggests that the model relies on pre-existing knowledge for familiar compounds, but struggles when required to reason about unseen combinations. Their analysis highlights the tension between reasoning and parroting in large models, making a case for the limits of model comprehension on noun compound tasks.
Further illustrating the limitations of context utilization, Cheng and Bhat (2024) show that pretrained LLMs often perform better in idiomatic expression reasoning when context is removed entirely. This counterintuitive result suggests that context can sometimes confuse models rather than help them. Moreover, Sun et al. (2021) found that LLMs rely on context only when the answer can be directly retrieved from it. These results suggest that reasoning performance almost always increases with context removal.
@@ 3 DICE: Dataset for Idiomatic Contrastive Evaluation
We introduce ’DICE’, a novel dataset for controlled evaluation of potentially idiomatic expressions balanced for sense. An overview of the curation process adopted for DICE is presented in Figure 1.
Figure 1: A visual illustration of our dataset curation process. We obtain a list of idioms using existing datasets, which GPT-4 is then prompted to generate sentences where the idiom is used literally. Expert then annotate the sentences to select the literal annotation in the literal subset, and that the figurative subset consists of the examples from existing datasets.
@@
Table 1: This table presents the characteristics of the subset of data used in our experiments and analysis. We have also gathered additional data, and a summary of that subset is available in Appendix B.
Expression Selection
To build our dataset, we compiled a comprehensive list of idiomatic expressions, focusing on both phrasal idioms and noun compound idioms. For phrasal idioms, we identified overlapping expressions from two established resources: MAGPIE (Haagsma et al., 2020) and SLIDE (Jochim et al., 2018). For noun compound idioms, we selected idiomatic expressions that were common to the NCTTI dataset (Garcia et al., 2021b) and AStitchInLanguageModels (Tayyar Madabushi et al., 2021). We specifically focused on idiomatic noun compounds, as the meaning of these expressions cannot be directly derived from the meanings of their individual words.
@@
Sentence Generation
We used GPT-4 (OpenAI et al., 2024) to generate sentences where a given idiom appears in a literal context, intentionally suppressing its figurative interpretation. The specific prompting setup we used for sentence generation is provided in Appendix C. To determine the best model for sentence generation, we piloted the process using three versions: GPT-4o, GPT-4, and GPT-3.5. The models were prompted to produce three different sentences, where the form of the idiom must be kept the same. We tested out the annotation process (detailed below) on a set of 20 idioms, and the Cohen’s kappa coefficient of inter-rater reliability for this pilot stage was 0.987. Overall, we found GPT-4 to perform the best at generating sentences where the figurative interpretation is suppressed. Our preference for GPT-4 aligns with the findings of Phelps et al. (2024), which demonstrate that off-the-shelf GPT-4 possesses relatively stronger idiomaticity knowledge as it performed consistently well across idiomaticity detection tasks compared to other off-the-shelf LLMs. In total, we obtained 2,349 sentences from GPT-4.
Expert Annotations
We recruited four experts with at least three years of university-level experience in Linguistics, compensated at a rate of £15/hour. The annotators were tasked with reviewing each sentence and deciding whether to accept it unconditionally, skip it, or reject it if the idiom’s figurative meaning could not be fully suppressed. In cases of rejection, annotators provided reasons such as ambiguity, figurative interpretation, change of form, or other issues. We provide information about this subset in Appendix B and also the briefing the annotators received in Appendix D. If an expression was skipped, a second annotator reviewed it to confirm if it should be discarded. Examples of sentences for each category are presented in Table 3. The reliability of the annotation process was high, as reflected by a Cohen’s kappa coefficient of 0.95, indicating strong inter-rater agreement.
@@ 3.1 Dataset Statistics
In total, our contrastive evaluation dataset consists of 2,066 sentences, featuring 402 expressions. A summary of the statistics of our dataset is presented in Table 1. In this paper we only use the subset whose literal leaning was confirmed by the annotators, however, we also release the excluded sentences to enable further analysis. However, we release the complete dataset, with an additional subset that can serve as a resource for additional analyses and for creating even more challenging datasets.
@@ 4 How Well Do LLMs Use Context for Idiomaticity?
4.1 Experimental Setup
Models
We evalute 13 models on the task of idiomatic sense disambiguation. These models are: GPT-4o [^2] https://platform.openai.com/docs/models/gpt-4o, GPT-3.5-Turbo Brown et al. (2020), FLAN-T5 models in the XXL, XL, Large, Small sizes (Chung et al., 2023), Llama 2 (7B, 13B, 70B) (Touvron et al., 2023), and Llama 3 models (incl. Llama 3.1 (405B, 8B, 70B) (Dubey et al., 2024). Additionally, we evaluated GPT-4 (OpenAI et al., 2024), which was used to generate the sentences. The prompts and computational resources used for the experiments are reported in the Appendix E.
To ensure a comprehensive and dependable assessment, each model was given three distinct yet semantically equivalent prompt versions. We assessed the models under both zero-shot and few-shot conditions.
Evaluation
To thoroughly evaluate the models’ performance, we employed three distinct evaluation settings:
@@ 4.1 Experimental Setup
    Accuracy: This setting includes two sub-evaluations: (1) Figurative Accuracy: We computed the accuracy of each model in correctly identifying the figurative uses of expressions within the figurative subset. (2) Literal Accuracy: We assessed the accuracy of the models in correctly identifying the literal uses of expressions within the literal subset. These evaluations measure the models’ ability to recognize idiomatic and literal meanings based on context.
@@
    Lenient Consistency: This metric rewards the model for consistently classifying all instances of an expression as either literal or figurative within a specific type. For example, if the model correctly classifies all literal variations of a given expression as “literal”, it earns a point for that expression. Similarly, the model is rewarded if it correctly identifies all figurative instances of an expression as “idiomatic”.
[…]
    $\frac{\sum_{x}\mathbf{1}\left(\forall i,\text{Pred}(x_{i}^{Lit})=\text{Lit})+\mathbf{1}(\forall i,\text{Pred}(x_{i}^{Fig})=\text{Fig}\right)}{2*\text{Number of unique expressions}}$ (1)
    where $x$ is a unique expression in the DICE, and $\mathbf{1}(\cdot)$ is the indicator function.
  • $\bullet$
    Strict Consistency: This is the most stringent evaluation. The model had to correctly identify all variations of an expression in both figurative and literal contexts to be rewarded. This setting assumes that a truly understanding model should correctly classify an idiom regardless of its context.
[…]
    $\frac{\sum_{x\in\mathcal{X}}\mathbf{1}\left(\forall i,\text{Prediction}(x_{i})=\text{True Label}(x_{i})\right)}{\text{Number of unique expressions}}$ (2)
@@
Table 2: Results $\pm$ 1 std (over 3 different sets of prompts) in zeroshot (top) and oneshot (bottom) conditions. Bold values represent the best performance for overall metric (incl.Strict Consistency) across models. GPT-4 is separated from the rest, as this model generated the evaluation sentences.

### v2

@@ Abstract
Human processing of idioms heavily depends on interpreting the surrounding context in which they appear. While large language models (LLMs) have achieved impressive performance on idiomaticity detection benchmarks, this success may be driven by reasoning shortcuts present in existing datasets. To address this, we introduce a novel, controlled contrastive dataset (DICE) specifically designed to assess whether LLMs can effectively leverage context to disambiguate idiomatic meanings. Furthermore, we investigate the influence of collocational frequency and sentence probability—proxies for human processing known to affect idiom resolution—on model performance. Our results show that LLMs frequently fail to resolve idiomaticity when it depends on contextual understanding, and they perform better on sentences deemed more likely by the model. Additionally, idiom frequency influences performance but does not guarantee accurate interpretation. Our findings emphasize the limitations of current models in grasping contextual meaning and highlight the need for more context-sensitive evaluation.
https://github.com/mi-m1/dice
@@
Maggie Mi1 Aline Villavicencio1,2,3 Nafise Sadat Moosavi1 1University of Sheffield 2University of Exeter 3Alan Turing Institute zmi1@sheffield.ac.uk a.villavicencio@exeter.ac.uk n.s.moosavi@sheffield.ac.uk
@@ 1 Introduction
Idiomatic expressions (IEs) are strange birds whose meaning may not be straightforwardly related to the meaning of the component words in isolation. For example, proficient English speakers understand “kick the bucket” not as “striking a metal container”, but as “to die”. Estimates suggest that there are 25,000 fixed expressions in English alone (Wèinreich, 1969), and a similar estimate is quoted for French (Gross, 1982). Notably, this figure is comparable to the order of magnitude of individual words in the lexicon (Jackendoff, 1997). This suggests that idioms are not mere linguistic curiosities but fundamental components of language.
Additionally, some of these expressions are ambiguous, “potentially idiomatic expressions" (PIEs) that can be interpreted either non-compositionally (figuratively or idiomatically) or compositionally (literally), depending on the context in which they appear. Accurately identifying the meaning of a PIE often depends on its context and is essential for numerous downstream applications, such as machine translation (Dankers et al., 2022; Barreiro et al., 2013; Salton et al., 2014; Fadaee et al., 2018), sentiment analysis (Williams et al., 2015; Liu et al., 2017), and automatic spelling correction Horbach et al. (2016). Beyond these applications, understanding idiomatic expressions in context is also crucial to grasp the underlying meaning of the text.
@@
Consequently, this allows models to exploit surface-level differences, such as changes in grammatical structure or insertions, as shortcuts for determining literalness, rather than relying on their understanding of idiomaticity in context as the task intends. This undermines the evaluation’s effectiveness by encouraging models to use superficial cues instead of deeper contextual comprehension.
In this paper, we address this gap by introducing DICE (Dataset for Idiomatic Contrastive Evaluation), a benchmark designed to evaluate the ability of LLMs to interpret idiomatic expressions in context. We focus on idioms, as these figurative expressions may serve as key indicators of a model’s linguistic understanding. Given the possibility of a predominant sense, DICE presents idiomatic expressions in both literal and figurative contexts and challenges models to rely on context for the correct interpretation. This contrastive evaluation forces models to distinguish between meanings based solely on context, preventing them from relying on memorized idioms. We hypothesize that if models do not depend on memorization, they will perform equally well in both senses of a PIE.
Moreover, in human language processing, factors such as frequency and familiarity (Swinney and Cutler, 1979; Nippold and Rudzinski, 1993), alongside context (Estill and Kemper, 1982; Gibbs Jr et al., 1989; Gibbs and Nayak, 1989), are known to influence idiom comprehension and processing speed (Tabossi et al., 2009). We investigate whether similar effects hold for language models. Specifically, beyond contextual understanding, we examine the influence of a languageintrinsic feature (expression frequency) and a modelintrinsic feature (sentence likelihood) on model performance.
@@
(1) We present DICE, a comprehensive and robust evaluation dataset, containing PIEs that occur in the same grammatical form across both figurative and literal contexts. (2) Through fine-grained evaluations, we find that models struggle to use context for idiomaticity processing. (3) Based on frequency estimates, we find that frequency is not a “free lunch”: whilst highly frequent idioms may be more likely to be disambiguated correctly, there is a trade-off in model performance between literal and figurative settings. (4) For models that demonstrate some capacity for idiomatic understanding, we observe a strong relationship between the likelihood of the contextual sentence and idiomaticity detection performance.
@@ 2 Related Works
Idiomaticity Detection.
The task of idiomaticity sense disambiguation (ISD), or idiomaticity detection, involves evaluating whether an expression is used literally or figuratively in a sentence (Liu and Hwa, 2018; Salehi et al., 2014; Senaldi et al., 2016; Gharbieh et al., 2016). This task is typically framed as binary classification. While large language models have achieved strong performance on existing ISD benchmarks (Phelps et al., 2024; Zeng and Bhat, 2021), it remains unclear whether this reflects true contextual understanding or reliance on memorized surface forms (Garcia et al., 2021b). Given the crucial role of context in resolving idiomaticity, it is essential to evaluate whether models are genuinely interpreting surrounding text or simply exploiting distributional cues.
Existing Datasets.
@@
The biggest dataset for idiomatic sense disambiguation is MAGPIE (Haagsma et al., 2020). MAGPIE contains a total of 56,622 PIE instances, across 1,756 idioms. However, a large amount of deviation of the form of an expression was allowed when MAGPIE was curated. As a result, most of the literal uses of PIEs involve modifications to the form of the expression (similar to the example in § 1). Other large datasets targeting various types of IEs have been released: The VNC-Tokens dataset focusing on verb-noun combinations (Cook et al., 2008), IDIX on verb-noun phrase or verb-prepositional phrase expressions (Sporleder et al., 2010). SemEval-2013 has unrestricted expressions (Korkontzelos et al., 2013), AStitchInLanguageModels contains noun compounds (Tayyar Madabushi et al., 2021) and more recently, IdioTS contains a mixture of expressions changed and unchanged to support the literal meaning (De Luca Fornaciari et al., 2024).
@@
To address these problems: (1) we propose a novel evaluation set (DICE), which strictly controls the form of idiomatic expressions. This design eliminates the possibility that models rely on grammatical variations for idiomaticity disambiguation. By maintaining consistent forms across literal and figurative contexts, DICE ensures that the challenge lies in understanding contextual nuances, thereby providing a more accurate assessment of a model’s idiomatic comprehension. (2) Existing datasets typically focus on a single type of expression. We address this limitation by including both phrasal idioms and noun compounds to provide broader coverage of idiomatic phenomena.
Contrastive Evaluation.
This evaluation paradigm involves comparing model outputs on carefully constructed input pairs that differ in minimal, controlled ways, typically along a dimension relevant to a specific linguistic phenomenon (Prasad et al., 2019; Garcia et al., 2021a). Such comparisons have been found to be particularly effective at isolating specific linguistic competencies or revealing systematic weaknesses in generalization and robustness (Linzen et al., 2016; Sennrich, 2017; Robertson, 2019). Our work adopts this approach to assess contextual comprehension in idiomaticity detection. Specifically, DICE presents potentially idiomatic expressions in both literal and figurative uses while holding surface form constant. This setup requires models to rely entirely on surrounding context to disambiguate meaning, eliminating the possibility of exploiting shallow lexical or syntactic cues.
Memorization and Context.
@@
Pretrained language models tend to handle IEs mainly through memorization of stored expressions and token distributions rather than reasoning about the meaning in context. This reliance on memorization becomes evident when models face novel noun compounds. For instance, Li et al. (2022) found that while GPT-3’s interpretations of novel compounds closely matched human responses, its performance faltered when tasked with interpreting nonsensical strings, which suggests limited contextual flexibility.
Similarly, Coil and Shwartz (2023) explored noun compound interpretation using GPT-3 and found that, although the model performs well with common noun compounds, its performance drops significantly with novel compounds. This suggests that the model relies on pre-existing knowledge for familiar compounds, but struggles when required to reason about unseen combinations. Supporting this view, Cheng and Bhat (2024) found that models sometimes perform better on idiomatic reasoning tasks when contextual information is removed, implying that context may occasionally mislead rather than assist the model. Sun et al. (2021) further showed that LLMs tend to use context only when the relevant information is explicitly present. These findings collectively raise questions about whether current models truly engage in contextual reasoning or primarily depend on surface-level cues.
Figure 1: Overview of the DICE curation process. A list of idioms is extracted from existing datasets, and GPT-4 is prompted to generate sentences using each idiom in a literal context. Experts annotate these to form the literal subset. Figurative examples are drawn directly from existing datasets.
@@ 3 DICE: Dataset for Idiomatic Contrastive Evaluation
We now describe the construction of DICE. Figure 1 shows an overview of the curation process.
@@
Table 1: Summary statistics of the DICE dataset.
Expression Selection.
To build our dataset, we compiled a comprehensive list of idiomatic expressions, focusing on both phrasal idioms and noun compound idioms. For phrasal idioms, we identified overlapping expressions from two established resources: MAGPIE (Haagsma et al., 2020) and SLIDE (Jochim et al., 2018). For noun compound idioms, we selected idiomatic expressions that appeared in both the NCTTI dataset (Garcia et al., 2021b) and AStitchInLanguageModels (Tayyar Madabushi et al., 2021). We specifically focused on idiomatic noun compounds, as the meaning of these expressions cannot be directly derived from the meanings of their individual words.
@@
Sentence Generation.
We used GPT-4 (OpenAI et al., 2024) to generate sentences where a given idiom appears in a literal context, intentionally suppressing its figurative interpretation. The specific prompting setup we used for sentence generation is provided in Appendix C. To determine the best model for sentence generation, we piloted the process using three versions: GPT-4o, GPT-4, and GPT-3.5. The models were prompted to produce three different sentences, where the form of the idiom must be kept the same. Overall, we found GPT-4 to perform the best in generating sentences where the figurative interpretation is suppressed. Our preference for GPT-4 aligns with the findings of Phelps et al. (2024), which demonstrate that off-the-shelf GPT-4 possesses relatively stronger idiomaticity knowledge as it performed consistently well across idiomaticity detection tasks compared to other off-the-shelf LLMs. In total, we obtained 2,349 sentences from GPT-4.
Expert Annotations.
We recruited four experts with at least three years of university-level experience in Linguistics, compensated at a rate of £15/hour. The annotators reviewed each sentence and decided whether to accept it unconditionally, skip it, or reject it if the idiom’s figurative meaning could not be fully suppressed. In cases of rejection, annotators provided reasons such as ambiguity, figurative interpretation, change of form, or other issues. We provide information about this subset in Appendix B along with the briefing given to annotators in Appendix D. If an expression was skipped, a second annotator reviewed it to confirm if it should be discarded. Examples of sentences for each category are presented in Table 3. The inter-rater agreement was high, as indicated by a Cohen’s kappa coefficient of 0.95.
@@ 3.1 Dataset Statistics
In total, DICE consists of 2,066 sentences, featuring 402 expressions. A summary of its statistics is presented in Table 1. In this paper, we only use the subset whose literal meaning was confirmed by the annotators. However, we release the complete dataset, along with an additional subset that can serve as a resource for additional analyses or for creating even more challenging evaluation settings (see Appendix B).
@@ 4 How Well Do LLMs Use Context for Idiomaticity?
4.1 Experimental Setup.
Models.
We evalute 13 models on the task of idiomatic sense disambiguation. These models are: GPT-4o, GPT-3.5-Turbo (Brown et al., 2020), FLAN-T5 models in the XXL (11B), XL (3B), Large (780M), Small (80M) sizes (Chung et al., 2023), Llama 2 (7B, 13B, 70B) (Touvron et al., 2023), and Llama 3 models (incl. Llama 3.1 (405B, 8B, 70B) (Dubey et al., 2024). Additionally, we evaluated GPT-4 (OpenAI et al., 2024), which was used to generate the sentences. The computational resources used are reported in Appendix E.
Prompts.
We used three prompt variations (Table 5 in § E.2) that are semantically equivalent but differ in surface form. This allows us to evaluate the robustness of model predictions to prompt phrasing. We report the mean and standard deviation across these prompts to assess performance stability under variation. We also run experiments in a few-shot setting, where an annotated example (shown in the middle part of Table 5) is prepended to each prompt.
Evaluation.
To thoroughly evaluate the models’ performance, we employed three distinct evaluation settings.
Accuracy includes two sub evaluations:
@@
    Figurative Accuracy: We compute the accuracy of each model in correctly identifying the figurative uses of expressions within the figurative subset. Let $F$ be the set of all figurative instances, $y_{i}$ and $\hat{y}_{i}$ denote the true and predicted labels, respectively.
    $\text{Accuracy}_{\text{fig}}=\frac{1}{|F|}\sum_{i\in F}\mathds{1}(\hat{y}_{i}=y_{i})$ (1)
    where
    $\mathds{1}(\text{condition})=\begin{cases}1&\text{if condition is true}\\ 0&\text{otherwise}\end{cases}$
@@
    Literal Accuracy: We assessed the accuracy of the models in correctly identifying the literal uses of expressions within the literal subset. These evaluations measure the models’ ability to recognize idiomatic and literal meanings based on context. Let $L$ be the set of all literal instances.
$\text{Accuracy}_{\text{lit}}=\frac{1}{|L|}\sum_{i\in L}\mathds{1}(\hat{y}_{i}=y_{i})$ (2)
@@
Table 2: Results are reported as mean $\pm$ standard deviation over 3 different prompt variants, under zero-shot (top) and one-shot (bottom) conditions. GPT-4 is shown separately, as it generated the evaluation sentences.
Lenient Consistency: This metric rewards the model for consistently classifying all instances of an expression as either literal or figurative within a specific type. For example, if the model correctly classifies all literal variations of a given expression as “literal”, it earns a point for that expression. Similarly, the model is rewarded if it correctly identifies all figurative instances of an expression as “idiomatic”. Let $E$ be the set of all expressions. Let $L_{e}$ and $F_{e}$ be the literal and figurative instances of expression $e$, respectively.
$C_{e}^{\text{lit}}=\begin{cases}1&\text{if }\forall i\in L_{e},\hat{y}_{i}=y_{i}\\ 0&\text{otherwise}\end{cases}$ (3)
$C_{e}^{\text{fig}}=\begin{cases}1&\text{if }\forall i\in F_{e},\hat{y}_{i}=y_{i}\\ 0&\text{otherwise}\end{cases}$ (4)
Lenient $\text{Lenient Consistency}=\frac{1}{2|E|}\sum_{e\in E}\left(C_{e}^{\text{lit}}+C_{e}^{\text{fig}}\right)$ (5)
Strict Consistency: This is the most stringent evaluation. The model had to correctly identify all variations of an expression in both figurative and literal contexts to be rewarded. This setting assumes that a truly understanding model should correctly classify an idiom regardless of its context. Let $V_{e}$ be the set of all variations of expression $e$.
$S_{e}=\begin{cases}1&\text{if }\forall i\in V_{e},\hat{y}_{i}=y_{i}\\ 0&\text{otherwise}\end{cases}$ (6)
Strict $\text{Strict Consistency}=\frac{1}{|E|}\sum_{e\in E}S_{e}$ (7)